# A Bayes Factor Framework for Unified Parameter Estimation and Hypothesis Testing

### v3

@@ 4.4 Logistic regression
Figure 7 shows the SCs related to a logistic regression analysis of the data including an intercept term and a main effect for each covariate. Each SC relates to the exponentiated regression coefficient, which can be interpreted as the multiplicative change of odds of street dealer presence when increasing the variable by one unit while keeping all other variables fixed. An improper flat prior was assigned to the intercept under both the null and the alternative, while independent $\operatorname{N}(0,1/2)$ priors were assigned to the coefficients under the alternative. This prior represents a weakly informative alternative postulating that the median odds ratio is $1$ and that $95\%$ of odds ratios are in between $1/4$ and $4$, representing a vague but realistic range of odds ratios in observational data (Greenland, , 2006). For the analysis of each coefficient, all other coefficients were considered as nuisance parameters with the same priors assigned to them under the null as under the alternative.
SCs were computed in three ways: i) By first computing the marginal posterior distribution for each coefficient from kernel smoothing of $0001^{\prime}000^{\prime}$ 000 Markov chain Monte Carlo (MCMC) samples (solid orange lines) computed with Stan (Carpenter et al., , 2017) and then computing the SC via the Savage-Dickey density ratio as explained in Section 3.4. ii) By computing the marginal posterior with integrated nested Laplace approximation (INLA; short-dashed blue lines) via INLA (Rue et al., , 2009) and then computing the SC via the Savage-Dickey density. iii) By estimating the parameters of the logistic model first with maximum likelihood, and then using each estimated coefficient and its standard error for a univariate normal analysis as explained in Section 2.3, ignoring the nuisance parameters (long-dashed green lines). As a result, the MEEs from the univariate normal analysis correspond to maximum likelihood estimates, while the MEEs from the MCMC and INLA analyses correspond to integrated maximum likelihood estimates.
@@
Concering the effect of the intervention, the SC (based on MCMC) for the variable ‘After intervention (top left panel) has its mode at $\widehat{\text{OR}}_{\mathrm{ME}}=1/1.8$ indicating a negative association between the text-based intervention and the presence of street dealers, yet this parameter value receives only moderate support over the alternative ($k_{\mathrm{ME}}=3.5$) and the corresponding $k=1$ support interval spans the range from ($\text{OR}=1/2.8$) up to hardly any association ($\text{OR}=1/1.1$). This suggests some negative association between the intervention and the presence of street dealers, although the extent of the association being relatively unclear. Moreover, the absence of a control condition makes it difficult to infer whether or not this association truly represents a causal effect.
@@ 5 Discussion
Like the likelihoodist and Neyman-Pearson paradigms of statistical inference, SC inference requires the formulation of alternative hypotheses. For this reason, SCs are particularly valuable in contexts where prior data or theories are available to formulate alternative hypotheses. For example, SCs (under the name of ‘K ratio‘) have been used by the large-scale NANOGrav collaboration to quantify the evidence for new physics theories against the established Standard Model (Afzal et al., , 2023). In cases where there are no clear alternative hypotheses, data analysts may use SCs based on ‘weakly informative‘ (Gelman, , 2009) or ‘default‘ prior distributions (e.g., unit-information priors, see Kass and Wasserman, , 1995), but should acknowledge this limitation and report sensitivity analyses (e.g., SCs for different prior distributions). Another possibility is to base SC inference on Bayes factor bounds (Berger and Sellke, , 1987; Sellke et al., , 2001; Held and Ott, , 2018), which give a bound on the maximum evidence against parameter values, but at the cost of losing the ability to quantify evidence in favour of parameter values (Pawel et al., , 2024).
@@
Bayesian, likelihoodist, or predictive reasoning may all motivate the Bayes factor as a natural tool for quantifying the relative evidence or support of competing hypotheses. Nevertheless, neither the Bayes factor nor any other measure of statistical evidence is infallible or suitable for all purposes. For example, Bayes factors by construction do not take into account the prior probabilities of their contrasted hypotheses, so they may indicate strong support for a hypothesis even though this hypothesis would still remain unlikely when combined with its prior probability (Lavine and Schervish, , 1999; Good, , 2001). Any type of statistical inference can lead to distorted scientific inferences if used in a bright-line fashion without consideration of contextual factors (Goodman, , 2016; Greenland, , 2023). We believe that SCs are useful in this regard because they shift the focus from finding evidence against a single null hypothesis to making gradual and quantitative inferences.
@@ Acknowledgments
We thank Ejbye-Ernst et al., (2023) and Bartoš et al., (2023) for openly sharing their data. We thank Riko Kelter, Andrew Fowlie, Eric-Jan Wagenmakers, František Bartoš, Leonhard Held, Małgorzata Roos, and Sander Greenland for valuable comments on drafts of the manuscript. The acknowledgment of these individuals does not imply their endorsement of the paper.

### v4

@@ 4.4 Logistic regression
Figure 7 shows the SCs related to a logistic regression analysis of the data including an intercept term and a main effect for each covariate. Each SC relates to the exponentiated regression coefficient, which can be interpreted as the multiplicative change of odds of street dealer presence when increasing the variable by one unit while keeping all other variables fixed. An improper flat prior was assigned to the intercept under both the null and the alternative, while independent $\operatorname{N}(0,1/2)$ priors were assigned to the coefficients under the alternative. This prior represents a weakly informative alternative postulating that the median odds ratio is $1$ and that $95\%$ of odds ratios are in between $1/4$ and $4$, representing a vague but realistic range of odds ratios in observational data (Greenland, , 2006). For the analysis of each coefficient, all other coefficients were considered as nuisance parameters with the same priors assigned to them under the null as under the alternative. Note that even though we only look at the SC of a single parameter, this approach does not ignore the nuisance parameters but averages the marginal likelihood over their prior, thereby incorporating nuisance parameter uncertainty.
SCs were computed in three ways: i) By first computing the marginal posterior distribution for each coefficient from kernel smoothing of $0001^{\prime}000^{\prime}$ 000 Markov chain Monte Carlo (MCMC) samples (solid orange lines) computed with Stan (Carpenter et al., , 2017) and then computing the SC via the Savage-Dickey density ratio as explained in Section 3.4. ii) By computing the marginal posterior with integrated nested Laplace approximation (INLA; short-dashed blue lines) via INLA (Rue et al., , 2009) and then computing the SC via the Savage-Dickey density. iii) By estimating the parameters of the logistic model first with maximum likelihood, and then using each estimated coefficient and its standard error for a univariate normal analysis similar to Section 4.3, ignoring the nuisance parameters in the SC analysis (long-dashed green lines). As a result, the MEEs from the univariate normal analysis correspond to maximum likelihood estimates, while the MEEs from the MCMC and INLA analyses correspond to integrated maximum likelihood estimates.
@@
Concering the effect of the intervention, the SC (based on MCMC) for the variable ‘After intervention (top left panel) has its mode at $\widehat{\text{OR}}_{\mathrm{ME}}=1/1.8$ indicating a negative association between the text-based intervention and the presence of street dealers, yet this parameter value receives only moderate support over the alternative ($k_{\mathrm{ME}}=3.5$) and the corresponding $k=1$ support interval spans the range from a noticeable association ($\text{OR}=1/2.8$) up to hardly any association ($\text{OR}=1/1.1$). This suggests some negative association between the intervention and the presence of street dealers, although the extent of the association being relatively unclear.
@@ 5 Discussion
Like the likelihoodist and Neyman-Pearson paradigms of statistical inference, SC inference requires the formulation of alternative hypotheses. For this reason, SCs are particularly valuable in contexts where prior data or theories are available to formulate alternative hypotheses. For example, SCs (under the name of ‘K ratio‘) have been used by the large-scale NANOGrav collaboration to quantify the evidence for new physics theories against the established Standard Model (Afzal et al., , 2023). Similarly, our application of SCs to the coin flipping experiment from Bartoš et al., (2025) or the replication studies from Wagenmakers et al., (2016) was relatively straightforward as the prior distribution for the parameters under the alternative could be specified through the a priori specified hypotheses and the data from the original study, respectively. In cases where there are no clear alternative hypotheses, data analysts may use SCs based on ‘weakly informative’ (Gelman, , 2009) or ‘default’ prior distributions (e.g., unit-information priors, see Kass and Wasserman, , 1995), but should acknowledge this limitation and report sensitivity analyses (e.g., SCs for different prior distributions). Another possibility is to base SC inference on Bayes factor bounds (Berger and Sellke, , 1987; Sellke et al., , 2001; Held and Ott, , 2018), which give a bound on the maximum evidence against parameter values, but at the cost of losing the ability to quantify evidence in favour of parameter values (Pawel et al., , 2024).
We have focused on situations where null models are nested within the alternative model. However, SC inference could also be applied to more general model comparison where the null is not nested in the alternative. In this case, the interpretation of the SC is more intricate as the comparison does not only involve an additional parameter but an entirely different model. Moreover, computing the SC via the Savage-Dickey density ratio is not possible anymore.
@@
Bayesian, likelihoodist, or predictive reasoning may all motivate the Bayes factor as a natural tool for quantifying the relative evidence or support of competing hypotheses. Nevertheless, neither the Bayes factor nor any other measure of statistical evidence is infallible or suitable for all purposes. In fact, there is empirical evidence that researchers sometimes misuse or misinterpret Bayes factors. For example, Bayes factors are sometimes erronously interpreted as absolute rather than relative evidence, or as posterior odds rather than updating factors (Wong et al., , 2022; Tendeiro et al., , 2024). Bayes factors by construction do not take into account the prior probabilities of their contrasted hypotheses, so they may indicate strong support for a hypothesis even though this hypothesis would still remain unlikely when combined with its prior probability (Lavine and Schervish, , 1999; Good, , 2001). Any type of statistical inference can lead to distorted scientific inferences if used in a bright-line fashion without consideration of contextual factors (Goodman, , 2016; Greenland, , 2023). We believe that SCs are useful in this regard because they shift the focus from finding evidence against a single null hypothesis to making gradual and quantitative inferences.
@@ Acknowledgments
We thank Ejbye-Ernst et al., (2023) and Bartoš et al., (2025) for openly sharing their data. We thank Riko Kelter, Andrew Fowlie, Eric-Jan Wagenmakers, František Bartoš, Leonhard Held, Małgorzata Roos, and Sander Greenland for valuable comments on drafts of the manuscript. We thank Yang Liu, Jonathan Williams, and another anonymous reviewer for helpful and constructive comments that substantially improved paper. The acknowledgment of these individuals does not imply their endorsement of the paper.